# Vireo: Leveraging Depth and Language for Open-Vocabulary Domain-Generalized Semantic Segmentation

### v2

@@ 5 Conclusion
Limitations Vireo assumes a reliable RGB camera, but in rare cases (e.g., occlusion, glare, hardware failure), the stream can be lost, impairing perception. Future work will explore multi-source setups—automatically switching to lidar, radar, or event-based cameras when RGB fails—to keep segmentation robust in all conditions.
Appendix / Supplemental Material
Appendix A Open Source Dataset and Experiments
To advance research in real-world open-vocabulary and domain-generalized semantic segmentation, we introduce AustinScapes, a new large-scale street-view dataset collected from the city of Austin, Texas, USA. The dataset comprises 100,000 high-resolution street-level images, each accompanied by pixel-level semantic annotations. The annotation schema follows the Cityscapes 19-class labels, ensuring compatibility with existing benchmarks.
AustinScapes captures a wide variety of urban scenes under diverse lighting, architectural styles, and traffic conditions, making it well-suited for evaluating both fine-grained recognition and robustness across real-world complexities. We benchmark our method and previous state-of-the-art approaches on AustinScapes to assess generalization performance. Figure 8 shows the visualization results of DGSS and OVSS methods on the AustinScapes dataset. Experimental results show that our framework maintains superior segmentation quality, particularly in challenging object categories and under domain shifts, further validating its effectiveness on unseen urban domains.
Figure 8: Key zero-shot segmentation examples from existing DGSS and OVSS methods, and our Vireo, on the AustinScapes dataset.
Appendix B Related Works
This section provides a more comprehensive discussion of the research areas that are highly relevant to our work: Open-Vocabulary Semantic Segmentation (OVSS) and Domain-Generalized Semantic Segmentation (DGSS). By examining these areas, we aim to highlight our contributions in bridging the gap between DGSS and OVSS.
B.1 Open-Vocabulary Semantic Segmentation
Existing OVSS methods [24, 10, 30, 3] primarily leverage the rich semantic knowledge embedded in pre-trained vision-language models (VLMs) such as CLIP [54]. Early approaches often adopt two-stage pipelines. For example, ZSSeg [56] classifies class-agnostic masks using CLIP, and MaskCLIP [23] adapts CLIP for instance mask proposal classification. Towards end-to-end solutions, LSeg [22] generates pixel-level embeddings aligned with text features, and OVSeg [1] further enhances this by learning discriminative pixel embeddings using a class-agnostic decoder. OpenSeg [57] presents a unified, language-driven framework for open-vocabulary segmentation.
To reduce dependency on region proposals, FreeSeg [2] introduces a grouping-based strategy to cluster pixels. Generalized decoders such as X-Decoder [58] enable language-conditioned pixel-level predictions. GroupViT [59] explores hierarchical region grouping with textual concepts. SegCLIP [11], a task-specific adaptation, enhances spatial understanding from CLIP for dense prediction tasks. More recently, OpenSeeD [60] aims to unify multiple segmentation tasks under the open-vocabulary setting via a conditioned decoder, and Cat-seg [10] improves multimodal matching by jointly modeling image and text in the cost function. Additionally, [61] addresses evaluation issues by proposing a mask-matching-based framework to handle semantic ambiguity.
Despite these advances, challenges remain in transferring coarse-grained VLM knowledge to pixel-level predictions and in resolving fine-grained linguistic ambiguities. Our work builds upon this line of research by incorporating depth cues to enhance language-guided segmentation, particularly in disambiguating objects within complex scenes.
B.2 Domain-Generalized Semantic Segmentation
Previous methods primarily focus on learning feature representations that are both discriminative for semantic segmentation and invariant to domain-specific variations. A foundational strategy in DGSS is introduced by DANN [27], which employs adversarial training with a domain classifier. Li et al. [62] investigate generative approaches using GANs to model the joint distribution of images and labels, while RobustNet [28] mitigates domain-specific biases through instance-selective whitening. SAN-SAW [13] introduces semantic-aware normalization using SAN and SAW modules, and BlindNet [63] enhances robustness by disentangling style and content representations.
More recently, pre-trained VFMs have been effectively leveraged in DGSS frameworks [8, 6, 64, 5]. These methods utilize frozen VFM features as universal backbones and optimize learnable tokens to improve cross-domain performance.
Building upon these advancements, our work posits that integrating the semantic generality of language with the geometric stability of depth provides a compelling approach to achieving robust domain-generalized open-vocabulary segmentation.
Appendix C Expend Experiments and Analysis
In this section, we present per-class method comparisons, model details, and visual results that are excluded from the main manuscript owing to page constraints.
C.1 More Data Analysis and Model Details
From Table 7, our approach achieves a mIoU of 66.73%, significantly outperforming other OVSS approaches. Compared to the second-best SED (48.16% mIoU), this represents an improvement of 18.57 percentage points, demonstrating a substantial enhancement in cross-domain generalization. Moreover, our approach attains the best performance in 18 categories.
For small targets prone to missed detections under extreme conditions (such as traffic lights and pedestrians), our method achieves 62.31% and 72.94% (17.34% and 12.88% higher than SED), respectively. The results reflect the robustness of our method in complex scenarios.
Table 7: Performance comparison across OVSS methods on per-class IoU and mIoU (City.→BDD.). Top three results are highlighted as best, second, and third, respectively(%).
Method road side. build. wall fence pole light sign vege terr. sky pers. rider car truck bus train moto. bicy. mIoU
EBseg 89.72 42.47 81.24 25.11 31.24 45.95 40.14 44.32 80.57 17.67 93.20 50.52 24.96 83.97 35.60 62.83 0.00 39.60 40.02 40.59
FC-CLIP 93.63 57.06 79.53 18.63 38.78 48.86 24.17 48.45 83.89 45.27 83.26 69.16 55.44 88.55 44.13 75.35 0.00 66.74 60.34 51.41
CAT-Seg 92.08 50.94 79.41 29.08 38.85 13.88 25.88 28.88 79.26 44.20 91.11 45.51 30.07 82.74 40.65 64.85 0.00 43.23 36.18 44.28
SED 93.78 60.35 82.90 30.07 44.48 39.26 44.97 46.22 83.40 45.55 92.53 60.06 29.23 86.92 43.64 42.94 0.31 43.88 42.12 48.16
Ours 95.40 68.08 88.29 34.81 46.11 60.95 62.31 59.80 87.26 51.10 95.27 72.94 62.63 90.93 53.82 83.08 22.35 74.40 58.37 66.73
Table 8: Comparison of per-class IoU across OVSS methods on the DELIVER dataset. Top three results are highlighted as best, second, and third, respectively(%).
Class Cloud Fog Night Rain Sun
Building 73.71 81.43 81.63 53.96 81.64 52.58 64.43 59.10 50.79 62.79 56.24 77.20 74.99 58.13 75.90 57.34 72.25 79.73 60.24 71.22 85.55 83.17 81.47 49.24 84.39
Fence 39.24 18.22 25.89 28.31 29.54 37.52 14.10 19.50 25.53 21.56 21.20 7.07 10.65 12.99 8.94 38.06 11.77 16.88 24.36 20.86 37.10 10.67 23.77 22.65 28.26
Other 0.00 0.00 0.00 0.11 0.04 0.00 0.00 0.00 0.02 0.06 0.00 0.00 0.00 0.00 0.13 0.00 0.00 0.00 0.00 0.17 0.00 0.00 0.00 1.88 0.03
Pedestrian 68.86 39.43 21.84 4.21 16.64 73.61 36.72 14.02 2.96 36.21 56.44 31.99 9.00 8.78 5.93 66.21 33.65 19.85 3.28 17.14 64.61 30.83 13.25 3.37 23.59
Pole 32.17 19.50 15.62 6.74 31.35 31.91 21.96 17.90 5.82 35.65 24.20 20.46 14.19 6.51 33.20 32.85 26.89 20.69 6.69 34.53 37.35 22.33 19.39 4.20 34.08
RoadLine 0.00 5.87 2.30 4.46 2.95 0.00 5.44 2.28 3.94 1.30 0.00 5.09 1.94 3.59 2.68 0.00 5.63 2.11 4.01 2.38 0.00 4.80 2.09 3.96 3.02
Road 91.70 29.72 89.55 0.00 65.80 91.72 24.25 89.15 0.01 80.96 90.78 26.22 86.28 0.00 64.23 91.65 32.11 88.48 0.01 67.85 91.94 22.69 88.78 0.00 63.06
SideWalk 71.50 57.22 57.36 59.27 59.04 66.91 53.90 55.25 47.48 43.27 69.18 56.95 58.65 54.13 38.89 69.50 56.40 56.95 52.37 45.44 71.16 48.96 55.14 44.70 27.36
Vegetation 69.36 61.11 62.12 29.30 39.67 68.58 58.86 58.96 33.79 35.20 38.46 40.39 31.46 24.22 20.33 66.69 62.29 62.37 28.17 31.89 68.85 58.12 59.30 54.17 68.62
Cars 63.84 60.20 57.84 66.49 65.73 68.99 60.90 58.78 72.55 68.84 77.26 66.19 60.66 74.09 67.51 64.80 59.94 53.38 62.80 60.64 70.38 65.32 59.65 66.64 64.05
Wall 3.46 0.08 2.05 2.67 4.14 0.30 0.00 0.00 0.08 0.43 2.11 1.46 0.00 1.47 1.53 0.47 0.41 0.01 0.49 0.23 4.07 5.95 0.00 0.29 3.80
TrafficSign 22.00 26.99 25.02 30.59 23.55 32.28 36.64 24.98 10.11 25.88 20.83 35.63 33.87 12.18 26.76 13.01 28.84 26.61 14.26 20.98 32.60 41.36 42.29 29.98 32.87
Sky 86.90 92.22 94.09 26.28 59.90 67.51 82.47 82.87 18.12 41.63 43.22 79.51 56.92 21.98 7.81 65.98 77.80 92.65 11.00 13.61 96.37 95.20 95.22 23.70 97.22
Ground 0.00 0.01 0.00 0.00 0.01 0.00 0.00 0.00 0.00 0.00 0.00 0.05 0.00 0.00 0.00 0.00 0.00 0.07 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Bridge 0.00 0.00 0.15 0.00 0.00 0.00 0.00 0.11 0.00 0.00 0.00 0.00 0.04 0.13 0.00 0.00 0.00 0.01 0.03 0.00 0.00 0.00 0.29 0.00 0.00
RailTrack 0.06 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.21 0.00 0.00 0.00 0.00 0.06 0.00 0.00 0.00 0.25 0.00 0.00 0.00 0.00 0.00
GroundRail 0.00 12.80 0.00 15.07 0.00 0.00 8.32 0.03 11.33 0.00 0.00 11.20 0.00 6.31 0.00 0.00 6.93 0.00 8.85 0.00 0.00 6.72 0.00 11.54 0.00
TrafficLight 57.73 25.71 23.39 24.94 13.05 67.39 41.82 29.52 10.24 21.85 51.48 30.92 21.47 2.01 4.45 58.03 23.86 22.72 4.60 10.50 61.49 39.72 25.74 3.17 17.78
Static 0.00 0.07 0.00 0.12 0.00 0.00 0.01 0.01 0.00 0.22 0.00 0.01 0.00 0.00 0.00 0.00 0.00 0.01 0.00 0.00 0.00 0.00 0.04 0.02 0.00
Dynamic 0.00 0.00 0.00 0.01 0.36 0.00 0.00 0.00 0.00 0.06 0.00 0.00 0.00 0.00 0.53 0.00 0.00 0.00 0.01 0.18 0.00 0.00 0.00 0.00 0.78
Water 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.68 0.00 0.00 0.00 0.38 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Terrain 48.31 32.42 38.57 15.22 4.35 45.35 27.11 35.20 17.08 2.42 37.71 13.06 28.41 4.41 0.45 47.53 25.24 33.96 17.23 9.46 43.44 25.18 31.03 15.49 5.16
TwoWheeler 51.66 37.99 16.31 20.55 12.10 60.03 41.90 22.45 29.29 11.42 48.04 31.61 14.29 25.50 13.74 68.66 52.24 41.47 34.45 41.11 60.42 47.01 38.37 31.28 40.16
Bus 0.00 0.01 0.00 3.76 0.00 0.00 6.11 0.00 4.45 0.00 0.00 0.93 0.00 0.00 0.00 0.00 0.00 0.00 23.64 12.30 0.00 0.00 0.00 1.73 0.05
Truck 28.00 9.22 41.70 45.42 55.62 53.19 43.36 49.80 68.80 60.14 50.18 33.77 10.99 50.29 14.40 54.24 45.11 45.65 52.43 47.98 67.95 64.92 69.59 55.30 66.04
Table 9: Efficiency and Performance Comparison
Method Iter. Train (4 BS) Mem. FPS Infer. Perform.
SED 40k 6h59min 20.8G 1.73 33min 57.0%
CAT-Seg 40k 3h3min 9.7G 1.66 35min 48.9%
FC-CLIP 40k 5h34min 18.7G 1.45 40min 58.7%
EBSeg 40k 6h34min 22.1G 0.62 94min 49.7%
REIN 40k 12h6min 13.8G 1.88 31min 69.4%
Vireo 40k 15h2min 18.4G 1.49 39min 73.0%
As shown in Table 8, on the DELIVER dataset, our method achieves the highest mIoU across all five extreme weather conditions, outperforming the runner-up by an average of 8–11 percentage points. Under Sun and Cloud conditions, large classes achieve over 80% IoU. In more adverse conditions, our method still maintains IoUs above 30%. Even in the most challenging Night scenario, we outperform others by a margin of 7.10%. For fine-grained categories such as TwoWheeler, Pedestrian, TrafficSign, and TrafficLight, our method consistently outperforms competitors, demonstrating strong recognition of low-contrast and small-scale objects. Although rare categories exhibit lower absolute IoUs, we still rank among the top in most of these cases. Overall, our approach demonstrates superior performance across both frequent and challenging classes under diverse weather conditions.
Table 9 compares training time, memory usage, FPS, inference time, and final mIoU (averaged across ACDC, BDD100K, and Mapillary) for various OVSS methods and the baseline, all evaluated under a unified 40k-iteration schedule with a batch size of 8. While Vireo’s training time (15 hours 2 minutes) and inference time (43 minutes) are not the shortest, it achieves the highest mIoU of 73.0%, utilizing 18.4 GB of memory at 1.36 FPS, substantially outperforming all other approaches.
Importantly, we follow the previous configuration of intermediate DINOv2 feature layers ($\{\hat{f}^{V}_{l_{i}}\}_{i=1}^{4},l_{i}\in\{8,12,16,24\}$), which has been demonstrated in prior work to offer an optimal trade-off between accuracy and efficiency. This ensures a fair and consistent comparison across all methods in our experiments.
C.2 More Visualization Results
Figure 9 presents a visual comparison of segmentation results between DGSS methods and ours on the Cityscapes $\rightarrow$ ACDC target domain. The first row shows the original outputs, while the second and third rows display zoomed-in crops for closer inspection. In the original images, other models clearly miss regions on the left sidewalk. In the second-row crops, our method yields much finer and more complete boundaries around small objects such as traffic lights and signs, detecting nearly every instance. In the third-row crops, even distant vehicles and signals are accurately segmented. Moreover, Vireo maintains continuous, clean segmentation of roads and sidewalks in low-contrast conditions (nighttime, rain, snow), without breaks or noise artifacts. It preserves highly consistent segmentation across large background areas like sky, buildings, and vegetation, suppressing “speckle” noise. It precisely captures very fine structures such as curbs and road markings. And it produces coherent, complete contours for dynamic objects (pedestrians, cyclists, vehicles). Whether in clear daytime or foggy, rainy nighttime scenarios, Vireo demonstrates superior cross-scenario robustness.
Figure 10 presents a visual comparison of segmentation results between DGSS methods and ours on the GTA5 $\rightarrow$ Cityscapes + BDD + Mapillary target domains. Across these varying scene distributions, RobustNet, HGFormer, and REIN often misclassify or omit road markings, crosswalks, clusters of pedestrians, and parked vehicles. In contrast, our Vireo consistently preserves the complete contours of dynamic objects and slender structures (e.g., lampposts, traffic signals). It segments curbs and lane lines continuously without breaks and delineates building–sky boundaries more sharply. It renders vegetation and other background regions with uniform, noise-free color blocks. Even for distant groups of pedestrians or in highly cluttered street scenes, Vireo accurately separates adjacent classes. This further demonstrates its superior generalization and robustness to unseen domains.
Figure 10: Key segmentation examples of existing DGSS methods and Vireo under the GTA5→City., BDD. and Map. unseen target domains.
Figure 11 presents a comparison of segmentation results between various OVSS methods and Vireo on the Cityscapes $\rightarrow$ ACDC target domain. The first column shows the original input images, followed by CAT-Seg, EBSeg, FC-CLIP, SED, Vireo, and finally the ground truth. As observed, CAT-Seg and EBSeg often yield fragmented road and sidewalk segmentations with background noise under extreme conditions such as rain, snow, and nighttime. FC-CLIP partially recovers large objects but still fails to detect small targets like traffic lights and thin poles. SED shows some improvement in road continuity but struggles with distant small objects and background consistency. In contrast, Vireo consistently produces continuous, precise segmentations of roads, lane markings, pedestrians, vehicles, and traffic signals across all weather and lighting conditions. It maintains low-noise, highly uniform background regions. Its outputs most closely match the ground truth, further demonstrating Vireo’s superior performance and robustness in open-vocabulary, cross-domain scenarios.
Figure 11: Key segmentation examples of existing OVSS methods and Vireo under the Citys. $\rightarrow$ ACDC unseen target domains under Night, Snow, Rain, and Fog conditions.
Figure 12 compares OVSS methods and our Vireo on the Cityscapes $\rightarrow$ DELIVERtransfer setting, which includes sunny, rainy, cloudy, foggy, and nighttime conditions. FC-CLIP preserves mask edges but suffers from widespread class errors. CAT-Seg and EBSeg yield fragmented road and sidewalk masks with spurious background artifacts under adverse conditions. SED improves overall continuity but still missegments slender structures like poles and signs in low-contrast scenarios. In contrast, Vireo consistently produces clear, coherent masks for zebra crossings, lane markings, sidewalks, pedestrians, vehicles, and poles. It effectively suppresses noise in sky, building, and vegetation regions. The outputs most closely match the ground truth, demonstrating superior robustness and generalization in extreme, unseen domains.
As shown in Figure 13, adding CMPE and GTP transforms the attention from a diffuse pattern into a concentrated focus on critical regions such as roads, vehicles, and pedestrians. In the Affinity Map, bright clusters reveal the model’s ability to capture long-range dependencies within the same object, while dark boundaries emphasize clear separation between different semantic regions. Together, these results confirm our method’s robust handling of semantic coherence and pixel-level affinities across diverse weather and lighting conditions.
Figure 12: Key segmentation examples of existing OVSS methods and Vireo under the Citys. $\rightarrow$ DELIVER unseen target domains under Cloud, Fog, Night, Rain, and Sun conditions.
Figure 13: Visualizations of attention maps and affinity maps under different scenes, where CMPE denotes the Coarse Mask Prior Embedding and GTP represents the GeoText Prompts.

### v3

@@ 5 Conclusion
Limitations Vireo assumes a reliable RGB camera, but in rare cases (e.g., occlusion, glare, hardware failure), the stream can be lost, impairing perception. Future work will explore multi‐source setups—automatically switching to lidar, radar, or event‐based cameras when RGB fails—to keep segmentation robust in all conditions.
Acknowledgements
This work was supported in part by the Natural Science Foundation of Xiamen, China, under Grants 3502Z202373036 and 3502Z202371019; in part by the National Natural Science Foundation of China under Grant 42371457; in part by the Natural Science Foundation of Fujian Province, China, under Grants 2025J01345 and 2023J01803; by the 2025 University of Glasgow Early Career Reward for Excellence on “Developing Vision–Language Models Enhanced by Geospatial Intelligence” and the 2025 University of Glasgow Early Career Mobility Scheme (ECMS) Award; and by the Program of China Scholarship Council under Grant No. 202506380088.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract’s and introduction’s key claims map exactly onto our contributions and scope
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In the conclusion section of the paper we discuss the limitations of the paper.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The theory assumptions are in the methodology section.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The model realization and implementation details are provided in the experimental section.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The dataset is publicly available and the code has been open sourced.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The Deployment Details & Parameter Settings are provided in the experiment section.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The evaluation protocols of these segmentation datasets do not require report the error bars.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The computation resources and details are discussed in experiment section.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The authors have read the code of ethics. The experiments are all on public datasets without obeying the code of ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: At the end of the conclusion section, the broader impacts have been discussed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This work does not have such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All external code, data, and pre-trained models we build on are cited, their original licenses are named, and our use conforms to those terms.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We will release our code base with included readme files.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This work does not involve crowdsourcing nor research with human subject
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This work does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: No large language models were used as part of the research methodology
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.